OpenStudy (therazor69):

does (31/2)2=(91/4)?

OpenStudy (anonymous):

multiply 31/2 by 2

OpenStudy (therazor69):

Woops.does (31/2)^2=(91/4)?

OpenStudy (anonymous):

so now it's just (31/2)(31/2)

OpenStudy (anonymous):

It doesn't though.

OpenStudy (therazor69):

its not (91/4)

OpenStudy (anonymous):

nope

OpenStudy (therazor69):

thanks

OpenStudy (anonymous):

yw

OpenStudy (therazor69):

Why is it not (explanation)

OpenStudy (anonymous):

because \[(31/2)^2 => (31/2)(31/2) = 961/4 \neq 91/4\]

OpenStudy (anonymous):

Medal and close it :D

OpenStudy (therazor69):

My bad im new